OpenStudy (princesssleelee):

A car is purchased for $25,500. After each year, the resale value decreases by35%. What will the resale value be after 4years? Use the calculator provided and round your answer to the nearest dollar.

OpenStudy (word2):

Price after 1 Year: 25,500 * .35 = 8,925 25,500 - 8,925 = $16,575 ---------> Price after 1 year Price after 2 years: $16,575* .35 = 5801.25 $16,575- 5801.25 = $10,773.75 ---------> Price after 2 years Price after 3 years: $10,773.75* .35 = 3770.8125 10,773.75 - 3770.8125 = $7,002.9375 ----------> Price after 3 years Price after 4 years: $7,002.9375 * .35 = 2451.02812 $7002.9375 - 2451.02812= $4551.90938 ------> Price after 4 years, rounded to the nearest dollar = $4552

OpenStudy (princesssleelee):

Could you help me with another one @word2

OpenStudy (word2):

sure

OpenStudy (princesssleelee):

A principal of $4000 is invested at 5% interest, compounded annually. How much will the investment be worth after 9years? Use the calculator provided and round your answer to the nearest dollar.

OpenStudy (princesssleelee):

4000* 0.05= 200 4000-200=3800 -----> First year? is this how i do it... but a total of 9 times?

OpenStudy (word2):

yeah basically , but i'll shorten it for you . just do this and you will get your answer 4000 (0.05)^9

OpenStudy (princesssleelee):

$0 ?

OpenStudy (word2):

no , also for this your not subtracting . your adding key word "compounded annually"

OpenStudy (word2):

4000* 0.05= 200 4000+200=4200 - first year 4200*0.05 = 210 4200+210 = 4300 - second year

OpenStudy (princesssleelee):

$5000? @word2

OpenStudy (princesssleelee):

nevermind

OpenStudy (word2):

you got it ?

OpenStudy (princesssleelee):

4300*0.05=215 4300+215=4515---------- 3rd year 4515*0.05=225.75 4515+225.75=4740.75------------4th year 4740.75*0.05=237.0375 4740.75+237.0375=4977.7875------------5th year 4977.7875*0.05=248.889375 4977.7875+248.889375=5226.67688------6th year 5226.67688*0.05=261.333844 5226.67688+261.333844=5488.01072-----7th year 5488.01072*0.05=274.400536 5488.01072+274.400536=5762.41126-----8th year 5762.41126*0.05=288.120563 5762.41126+288.120563=6050.53182-----9th year Nearest Dollar ---- $6,051 @word2

OpenStudy (agent0smith):

@princesssleelee you could just use the compound interest formula...

OpenStudy (princesssleelee):

Im unsure of what that is, however, working it out, did i do it correctly?

OpenStudy (princesssleelee):

No, I just submitted my answer and it was incorrect

OpenStudy (agent0smith):

No way am i checking all that work lol, when there's a single formula for this. https://qrc.depaul.edu/StudyGuide2009/Notes/Savings%20Accounts/Compound%20Interest_files/image002.gif P is the initial amount r is the interest rate as a decimal n is number of times compounded per year t is number of years

OpenStudy (word2):

that way does takes a lot of work

OpenStudy (word2):

try using the formula

OpenStudy (princesssleelee):

6205

OpenStudy (princesssleelee):

A city has a population of 390,000 people. Suppose that each year the population grows by 8.5%. What will the population be after 5 years? so for this would I use the same formula, doing so gave me 30177456562.5

OpenStudy (princesssleelee):

And it was incorrect...... :/

OpenStudy (agent0smith):

Of course it's incorrect. Gotta pay attention to whether your answers seem reasonable. That number is larger than 30 billion. The population of the entire planet is 7 billion.

OpenStudy (princesssleelee):

I looked at that, but i do not know any other way to figure out the problem.

OpenStudy (agent0smith):

Show your work.

OpenStudy (princesssleelee):

What do you mean?

OpenStudy (agent0smith):

By show your work? I mean... show your work.

OpenStudy (princesssleelee):

I used a calculator.

OpenStudy (agent0smith):

Then you typed it in wrong. Show what you typed in.

OpenStudy (princesssleelee):

390000(1+8.5/1)^5

OpenStudy (agent0smith):

You missed something i wrote earlier "r is the interest rate as a decimal" I assume you know how to convert a percentage to a decimal.

OpenStudy (princesssleelee):

8.5

OpenStudy (agent0smith):

It's 8.5%, convert it to a decimal.

OpenStudy (princesssleelee):

0.085

OpenStudy (princesssleelee):

586426 :)

OpenStudy (agent0smith):

Looks reasonable.

OpenStudy (princesssleelee):

Thank you very much!!